OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim?

OpenStudy (anonymous):

15 years old..

OpenStudy (anonymous):

How? U see, here, it is given that dad = 4(jim) after 10 yearsv .. dad's age will be (dad+10) and jim will be (jim + 10) Put these values accordin to the situation given in the question .. i.e dad+10 = 2(jim+10) + 20 =. 4(jim) + 10 = 2(jim+10) + 20.. Hope you understood. :)